### v1

@@ Abstract
Given two points in the plane, and a set of “obstacles” given as curves through the plane with assigned weights, we consider the point-separation problem, which asks for the minimum-weight subset of the obstacles separating the two points. A few computational models for this problem have been previously studied. We give a unified approach to this problem in all models via a reduction to a particular shortest-path problem, and obtain improved running times in essentially all cases. In addition, we also give fine-grained lower bounds for many cases.
@@ 1 Introduction
Given points $s$ and $t$ in the plane, and a weighted set of obstacles $\mathcal{C}$ defined by simple closed curves (possibly also including their interiors), the $(s,t)$ point-separation problem asks for the minimum-weight subset $C$ of $\mathcal{C}$ such that any path from $s$ to $t$ intersects some obstacle in $C$. Equivalently, $s$ and $t$ are in different connected components of $\mathbb{R}^{2}\setminus(\cup_{\gamma\in C}\gamma)$. We say such a subset separates $s$ and $t$. An example of this problem can be found in fig. 1.
This is a natural problem that arises in various scenarios. As a toy application of this problem: Suppose every night your dog runs from his doghouse to your backpack to eat your homework, taking an unpredictable route (making use of windows and doors). You have noticed that your dog will forget about your homework if it smells a treat. You have a number of candidate locations to place treats, and you’d like to ensure your dog is distracted from your homework every day using the fewest treats possible. See fig. 2 for this example. Similar applications also arise when considering security (e.g., replacing the batteries in the fewest number of your dead security cameras to cover all possible paths from the entrance to your bank vault). Additionally, $(s,t)$ point-separation has found an application as a tool for constant-factor approximation of the well-studied APX-hard problem “barrier-resilience” [KumarLSS21].
@@
  • The Oracle-based Intersection Graph Model: Assume the existence of several $O(1)$-time oracles that would allow the computation of the intersection graph $G$ of $\mathcal{C}$ (the graph over vertex set $\mathcal{C}$ whose edges correspond to pairwise intersections), as well as some additional information for each edge related to $s$ and $t$ (detailed in section 2).
  • The Arrangement Model: Assume the arrangement of $\mathcal{C}$ is provided as input, in the form of a plane (multi-)graph, with the faces corresponding to $s$ and $t$ labelled. This is a more graph-theoretic formulation. See fig. 3 for an example of such an arrangement.
Figure 3: An arrangement representation of the instance in fig. 1.
@@ 1.1 Prior Work (Brief)
Most importantly for our methods, Kumar, Lokshtanov, Saurabh and Suri [KumarLSS21, Section 6] describe an algorithm that runs in polynomial time in the arrangement model. Their algorithm implicitly makes use of the homology cover (perhaps unintentionally), in the same sense that we will use it. In fact, their algorithm is some ways analogous to an algorithm of Chambers, Erickson, Fox, and Nayyeri [ChambersEFN23] for minimum-cuts (and maximum-flow) in surface graphs. These two algorithms are the main inspiration for our approach to the $(s,t)$ point-separation at a high-level, in all models.
@@ 2 Homology and Obstacles [Informal]
In this section, we will explain the main topological tools we will use to approach the $(s,t)$ point-separation problem. However, this will be an information section, not requiring prior knowledge of any aspect of topology. Rather, we will give an equivalent formulation of the important pieces using simpler tools from computational geometry. We give a more formal treatment in LABEL:sec:homology-formal Throughout this section, we will make (implicit) assumptions of general position and such, but LABEL:sec:homology-formal does not need these.
At a high-level, there are two main steps to the constructions of our approach. First, we will discuss what it means for a simple curve to separate $s$ and $t$, and what tools exist to classify such curves. Second, we will discuss what it means for a set of obstacles to separate $s$ and $t$. That is, when the union of the obstacles contains a simple curve separating $s$ and $t$¿
The simplest possible case of asking whether a simple curve separates $s$ and $t$ is characterized by the point-in-polygon problem, which asks: Given a point $p$ and a (simple) polygon $P$, is $p$ inside $P$? In fact, this is equivalent to asking if $P$ separates $p$ and the point at infinity. There is a folklore algorithm for this problem that takes any ray $r$ starting at $p$, and counts the number of intersections between $r$ and $P$. Then, $P$ contains $p$ if and only if the number of intersections is odd. In fact, essentially the same algorithm can be used to test whether or not $P$ separates two points $p_{1}$ and $p_{2}$. Rather than using a ray, take the segment $\overline{p_{1}p_{2}}$. Count the number of intersections between $\overline{p_{1}p_{2}}$ and $P$. The count is odd if and only if $P$ separates $p_{1}$ and $p_{2}$. Moreover, there is in fact nothing special about rays or line segments in this problem. Any (closed) path between $p_{1}$ and $p_{2}$ would cross $P$ the same number of times, modulo $2$. See fig. 4 for examples for these algorithms.
@@
All three of these algorithms are equivalent in a sense. In fact, there is a further generalization: Given two points $s$ and $t$ on the sphere, a simple and closed curve $C$ (not covering $s$ or $t$), and any $s-t$ path $\pi$, $C$ separates $s$ and $t$ if and only if $\pi$ crosses $C$ an odd number of times (regardless of the choice of $\pi$). Since the extended plane is homeomorphic to the sphere, a ray in the plane corresponds to a (simple) path in the sphere. See fig. 5 for an example.
@@
The problems we have discussed so far are all completely static, and the methods do not provide much structure for solving more difficult problems. One of the most common ways to extend the point-in-polygon problem is to fix the polygon (or curve) $P$, and aim to support fast queries of points. This problem is known as “point-location”, and it is well-studied in computational geometry [BergCKO08] However, we want a different sort of structure: We have a fixed pair of points $s$ and $t$, and we wish to classify the curves that separate them. Since we have fixed $s$ and $t$, we can also fix the path $\pi$ between them – in most cases, we will use the line segment $\overline{st}$. Then, the problem of classifying curves that separate $s$ and $t$ becomes the problem of classifying curves that cross $\overline{st}$ an odd number of times.
It will be helpful for demonstration to make some transformations to the space we work in. That is, we will perform a sequence of homeomorphisms. We start with the extended plane with the two marked points $s$ and $t$, and the line segment $\overline{st}$ (see fig. 9). No curves we will be classifying cross $s$ or $t$, so we may assume there are punctures at $s$ and $t$. We can expand these punctures with a homeomorphism (see fig. 9). Next, since we have the extended plane, we can make one of the punctures the outer face, obtaining an annulus (see fig. 9). In performing these steps, the line segment $\overline{st}$ becomes a path between the inner and outer boundaries of the annulus.
{subfigure}
[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption
Figure 6: $s$ and $t$ are punctures.
Figure 7: $s$ and $t$ are large punctures.
Figure 8: $s$ is the outer face, and the space is an annulus.
@@
We now present a method for constructing an important space called the homology cover111 We use the term “homology cover” to refer to the one-dimensional $\mathbb{Z}_{2}$-homology cover of the annulus. . Take the specified path between the two boundaries (see fig. 13) and slice it open (see fig. 13). Then, create a second copy of the sliced annulus, and glue them together in a way that matches up the orientations of the sliced ends (see fig. 13).
{subfigure}
[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption
Figure 10: An annulus with a specified path.
Figure 11: The annulus is sliced open at the specified path.
Figure 12: Two copies of the sliced annulus are glued together.
@@
Alternatively, an equivalent construction is to create two copies of the original space (whether that be the extended plane or the annulus), and use each side of the path from $s$ to $t$ as a (separate) “portal” between the two copies. A more general form of this “portal” idea has been studied in the form of “portalgons” [loffler_et_al, ophelders2024shortest], of which the homology cover is essentially a special case. See fig. 14 for an example of this construction.
@@
One important aspect of this construction is that it involves the connection of two identical copies of the annulus (i.e., the extended plane with punctures $s$ and $t$). In fig. 13, we also show how this construction transforms two curves – one separating the two boundaries, and one not separating them. The structure we will make use of to study curves separating $s$ and $t$ in the plane ultimately stems from an important set of facts:
@@ 2.1 Obstacles and the Homology Cover
In the arrangement model of the $(s,t)$ point-separation problem, we are given a plane graph $D$ (the arrangement) with specified faces $s$ and $t$ (can be obtained from points via point-location), and a set of obstacles $\sigma$, each given as a connected subgraph. The homology cover has a simple graph-theoretic interpretation in this framework: Take any (simple) dual-path $\pi$ of faces from $s$ to $t$ – this is will serve as the “portal”. Next, create two copies $D_{1},D_{2}$ of $D$, each with the faces $s$ and $t$ removed. We will create a combined graph $\overline{D}$ starting from the union of $D_{1}$ and $D_{2}$. For each edge $e=uv$ used in the dual-path $\pi$, delete $e$ from both $D_{1}$ and $D_{2}$ (inside $\overline{D}$). Denote the corresponding vertices of $u$ and $v$ in each of $D_{1}$ and $D_{2}$ as $u_{1},v_{1}$ and $u_{2},v_{2}$, respectively. Then create new edges $u_{1}v_{2}$ and $u_{2}v_{1}$ in $\overline{D}$. This form of the homology cover is visualized in fig. 15.
Figure 15: The homology cover of the arrangement given in fig. 3. The purple edges are the modified “crossing” edges.
With this particular arrangement structure Kumar, Lokshtanov, Saurabh and Suri [KumarLSS21, Section 6] built a graph using one vertex per obstacle-vertex incidence (in each copy of the plane graph). We will build a smaller graph of similar form to theirs. First, for each obstacle $C\in\sigma$ (which induces a subgraph of $D$), pick some arbitrary canonical point $x\in C$. Since this choice is arbitrary, it will sometimes be useful to assume it is a specific point – this will primarily be useful for specific obstacle types. Assume that $x$ is also a vertex in the subgraph of $D$ induced by $C$ (and if it is not, modify $D$ so that it is, with either a bisection or a new degree-$1$ vertex). Note that since $C$ is connected in $D$, every other vertex $x^{\prime}\in C$ is connected to $x$ by only edges in $C$. Given a plane graph $D$ with faces $s$ and $t$ a dual-path $\pi$, obstacles $\sigma$, and canonical points for each obstacle, the auxiliary graph $H$ is a bipartite graph constructed as follows:
@@
    • $v\in\gamma$ (that is, the point $v$ is covered by the obstacle $\gamma$ in the arrangement), and…
    • The canonical point $x$ of $\gamma$ has a path through the edges of $\gamma$ to $v$ whose intersection with the edges of $\pi$ is size $b_{1}+b_{2}$ modulo $2$. Equivalently, the projected connected component of $\gamma$ containing the canonical point copy $(\gamma,b_{2})$ also contains $v^{b_{1}}$.
The auxiliary graph is useful because the shortest-path between any two corresponding vertex copies or (separately) any two corresponding canonical point copies both correspond exactly to the solution to the $(s,t)$ point-separation problem. The high-level construction to prove this fact is as follows:
@@
Figure 16: The intersection graph (left) and the intersection graph in the homology cover (right) for the curves given in fig. 1. The purple edges are the “crossing” edges.
For the oracle model, the purpose of the oracles will be to construct the geometric intersection graph of the obstacle copies in the homology cover (which we will denote as $\overline{G}$). We will call $\overline{G}$ the intersection graph in the homology cover. See fig. 16 for an example. We will present two different constructions of this graph. They are equivalent, and each of them is quite simple, but they will serve two different purposes: The first will clarify which types of oracle queries are necessary, and provide an algorithmic construction. The second will instead build on the tools we have for the arrangement model, proving the correctness of another shortest-paths approach.
Let $\mathcal{C}$ denote the set of obstacles in the plane. We assume the canonical points for each obstacle are given (or implied) – they will be used by the oracle. We also assume some simple $s-t$ path $\pi$ is fixed – this will also be used by the oracle. As mentioned In most cases, $\pi$ will be $\overline{st}$. The first construction of the graph is as follows:
@@
Testing for this condition is actually the only type of oracle query needed (although we will also require support for the case when $\gamma_{1}=\gamma_{2}$ later). Cabello and Giannopoulos [CabelloG16] used a larger set of query types, but together they can be used to perform this type by carefully choosing the canonical points and fixing $\pi$ to $\overline{st}$, so our variant of the oracle model slightly more general (although practically equivalent).
@@
  • Therefore, the set of vertices visited by the shortest path from some $(\gamma,-)$ to $(\gamma,+)$ (over all possible $\gamma\in\mathcal{C}$) corresponds to the minimum-weight set of obstacles separating $s$ and $t$.

### v2

@@ Abstract
Given two points in the plane, and a set of “obstacles” given as curves through the plane with assigned weights, we consider the point-separation problem, which asks for a minimum-weight subset of the obstacles separating the two points. A few computational models for this problem have been previously studied. We give a unified approach to this problem in all models via a reduction to a particular shortest-path problem, and obtain improved running times in essentially all cases. In addition, we also give fine-grained lower bounds for many cases.
@@ 1 Introduction
Given points $s$ and $t$ in the plane, and a weighted set of obstacles $\mathcal{C}$ defined by simple closed curves (possibly also including their interiors), the $(s,t)$ point-separation problem asks for a minimum-weight subset $C$ of $\mathcal{C}$ such that any path from $s$ to $t$ intersects some obstacle in $C$. Equivalently, $s$ and $t$ are in different connected components of $\mathbb{R}^{2}\setminus(\cup_{\gamma\in C}\gamma)$. We say such a subset separates $s$ and $t$. An example of this problem can be found in Figure 1.
This is a natural problem that arises in various scenarios. As a toy application of this problem: Suppose every night your dog runs from his doghouse to your backpack to eat your homework, taking an unpredictable route (making use of windows and doors). You have noticed that your dog will forget about your homework if it smells a treat. You have a number of candidate locations to place treats, and you’d like to ensure your dog is distracted from your homework every day using the fewest treats possible. See Figure 2 for this example. Similar applications also arise when considering security (e.g., replacing the batteries in the fewest number of your dead security cameras to cover all possible paths from the entrance to your bank vault). Additionally, $(s,t)$ point-separation has found an application as a tool for constant-factor approximation of the well-studied APX-hard problem “barrier-resilience” [KumarLSS21].
@@
  • The Oracle-based Intersection Graph Model: Assume the existence of several $O(1)$-time oracles that would allow the computation of the intersection graph $G$ of $\mathcal{C}$ (the graph whose vertices are exactly the curves $\mathcal{C}$ and whose edges correspond to pairwise intersections of these curves), as well as some additional information for each edge related to $s$ and $t$ (detailed in Section 2).
  • The Arrangement Model: Assume the arrangement of $\mathcal{C}$ is provided as input, in the form of a plane (multi-)graph, with the faces corresponding to $s$ and $t$ labelled. This is a more graph-theoretic formulation. See Figure 3 for an example of such an arrangement.
Figure 3: An arrangement representation of the instance in Figure 1, as well as an arrangement of a smaller instance.
@@ 1.1 Prior Work (Brief)
Most importantly for our methods, Kumar, Lokshtanov, Saurabh and Suri [KumarLSS21, Section 6] describe an algorithm that runs in polynomial time in the arrangement model. Their algorithm implicitly makes use of the homology cover (perhaps unintentionally), in the same sense that we will use it. In fact, their algorithm is in some ways analogous to an algorithm of Chambers, Erickson, Fox, and Nayyeri [ChambersEFN23] for minimum-cuts (and maximum-flow) in surface graphs. These two algorithms are the main inspiration for our approach to the $(s,t)$ point-separation at a high-level, in all models.
@@ 2 Homology and Obstacles [Informal]
In this section, we will explain the main topological tools we will use to approach the $(s,t)$ point-separation problem. However, this will be an information section, not requiring prior knowledge of any aspect of topology. Rather, we will give an equivalent formulation of the important pieces using simpler tools from computational geometry. We give a more formal treatment in LABEL:sec:homology-formal. Throughout this section, we will make (implicit) assumptions of general position and such, but the deferred formal treatment does not need these.
At a high-level, there are two main steps to the constructions of our approach. First, we will discuss what it means for a simple curve to separate $s$ and $t$, and what tools exist to classify such curves. Second, we will discuss what it means for a set of obstacles to separate $s$ and $t$. That is, when does the union of the obstacles contain a simple curve separating $s$ and $t$?
The simplest possible case of asking whether a simple curve separates $s$ and $t$ is characterized by the point-in-polygon problem, which asks: Given a point $p$ and a (simple) polygon $P$, is $p$ inside $P$? In fact, this is equivalent to asking if $P$ separates $p$ and the point at infinity. There is a folklore algorithm for this problem that takes any ray $r$ starting at $p$, and counts the number of intersections between $r$ and $P$. Then, $P$ contains $p$ if and only if the number of intersections is odd. In fact, essentially the same algorithm can be used to test whether or not $P$ separates two points $p_{1}$ and $p_{2}$. Rather than using a ray, take the segment $\overline{p_{1}p_{2}}$. Count the number of intersections between $\overline{p_{1}p_{2}}$ and $P$. The count is odd if and only if $P$ separates $p_{1}$ and $p_{2}$. Moreover, there is in fact nothing special about rays or line segments in this problem. Any (closed) path between $p_{1}$ and $p_{2}$ would cross $P$ the same number of times, modulo $2$. See Figure 4 for examples for these algorithms.
@@
All three of these algorithms are equivalent in a sense. In fact, there is a further generalization: Given two points $s$ and $t$ on the sphere, a simple and closed curve $C$ (not covering $s$ or $t$), and any $s-t$ path $\pi$, $C$ separates $s$ and $t$ if and only if $\pi$ crosses $C$ an odd number of times (regardless of the choice of $\pi$). Since the extended plane is homeomorphic to the sphere, a ray in the plane corresponds to a (simple) path in the sphere. See Figure 5 for an example.
@@
The problems we have discussed so far are all completely static, and the methods do not provide much structure for solving more difficult problems. One of the most common ways to extend the point-in-polygon problem is to fix the polygon (or curve) $P$, and aim to support fast queries of points. This problem is known as “point-location”, and it is well-studied in computational geometry [BergCKO08]. However, we want a different sort of structure: We have a fixed pair of points $s$ and $t$, and we wish to classify the curves that separate them. Since we have fixed $s$ and $t$, we can also fix the path $\pi$ between them – in most cases, we will use the line segment $\overline{st}$. Then, the problem of classifying curves that separate $s$ and $t$ becomes the problem of classifying curves that cross $\overline{st}$ an odd number of times.
It will be helpful for demonstration to make some transformations to the space we work in. That is, we will perform a sequence of homeomorphisms. We start with the extended plane with the two marked points $s$ and $t$, and the line segment $\overline{st}$ (see Figure 6(a)). No curves we will be classifying cross $s$ or $t$, so we may assume there are punctures at $s$ and $t$. We can enlarge these punctures into holes with a homeomorphism. Next, since we have the extended plane, we can make one of the punctures the outer face, obtaining an annulus (see Figure 6(b) and Figure 6(c)). In performing these steps, the line segment $\overline{st}$ becomes a path between the inner and outer boundaries of the annulus.
(a) $s$ and $t$ are punctures, $\pi$ is a path between them.
(b) $s$ and $t$ are enlarged to become holes, and the space is an annulus.
(c) The two sides of $\pi$ can be coloured to aid later notation.
@@
We now present a method for constructing an important space called the homology cover111 We use the term “homology cover” to refer to the one-dimensional $\mathbb{Z}_{2}$-homology cover of the annulus. . Take the specified path $\pi$ between the two boundaries (see Figure 7(a)) and slice it open (see Figure 7(b)). Then, create a second copy of the sliced annulus, and glue them together in a way that matches up the orientations of the sliced ends (see Figure 7(c)).
(a) An annulus with a specified path $\pi$.
(b) The annulus is sliced open along the specified path $\pi$.
(c) Two copies of the sliced annulus are glued together.
@@
Alternatively, an equivalent construction is to create two copies of the original space (whether that be the extended plane or the annulus), and use each side of the path from $s$ to $t$ as a (separate) “portal” between the two copies. A more general form of this “portal” idea has been studied in the form of “portalgons” [loffler_et_al, ophelders2024shortest], of which the homology cover is essentially a special case. See Figure 8 for an example of this construction.
@@
One important aspect of this construction is that it involves the connection of two identical copies of the annulus (i.e., the extended plane with punctures $s$ and $t$). In Figure 7, we also show how this construction transforms two curves – one separating the two boundaries, and one not separating them. The structure we will make use of to study curves separating $s$ and $t$ in the plane ultimately stems from an important set of facts:
@@ 2.1 Obstacles and the Homology Cover
In the arrangement model of the $(s,t)$ point-separation problem, we are given a plane graph $D$ (the arrangement) with specified faces $s$ and $t$ (can be obtained from points via point-location), and a set of obstacles $\sigma$, each given as a connected subgraph. The homology cover has a simple graph-theoretic interpretation in this framework: Take any (simple) dual-path $\pi$ of faces from $s$ to $t$ – this is will serve as the “portal”. Next, create two copies $D_{1},D_{2}$ of $D$, each with the faces $s$ and $t$ removed. We will create a combined graph $\overline{D}$ starting from the union of $D_{1}$ and $D_{2}$. For each edge $e=uv$ used in the dual-path $\pi$, delete $e$ from both $D_{1}$ and $D_{2}$ (inside $\overline{D}$). Denote the corresponding vertices of $u$ and $v$ in each of $D_{1}$ and $D_{2}$ as $u_{1},v_{1}$ and $u_{2},v_{2}$, respectively. Then create new edges $u_{1}v_{2}$ and $u_{2}v_{1}$ in $\overline{D}$. This form of the homology cover is visualized in Figure 9.
(a) The arrangement with a specified path $\pi$.
(b) A simple separating cycle in the arrangement.
(c) The edges of the arrangement crossing $\pi$ are deleted.
(d) Two copies of are created, and the deleted edges are added as “crossing” edges (in purple).
(e) A path between two copies of the same original vertex that would project onto the simple separating cycle in the original space from Figure 9(b).
Figure 9: The construction of the homology cover for the second arrangement given in Figure 3.
With this particular arrangement structure Kumar, Lokshtanov, Saurabh and Suri [KumarLSS21, Section 6] built a graph using one vertex per obstacle-vertex incidence (in each copy of the plane graph). We will build a smaller graph of similar form to theirs. First, for each obstacle $C\in\sigma$ (which induces a subgraph of $D$), pick some arbitrary canonical point $x\in C$. Since this choice is arbitrary, it will sometimes be useful to assume it is a specific point – this will primarily be useful for specific obstacle types. Assume that $x$ is also a vertex in the subgraph of $D$ induced by $C$ (and if it is not, modify $D$ so that it is, with either a subdivision or a new degree-$1$ vertex). Note that since $C$ is connected in $D$, every other vertex $x^{\prime}\in C$ is connected to $x$ by only edges in $C$. Given a plane graph $D$ with faces $s$ and $t$, a dual-path $\pi$, obstacles $\sigma$, and canonical points for each obstacle, the auxiliary graph $H$ is a bipartite graph constructed as follows:
@@
  • The values -- and $+$ in each of the above are indicator bits, and can be stored as 00 and $1$, respectively.
@@
    • $v\in\gamma$ (that is, the point $v$ belongs to the obstacle $\gamma$ in the arrangement), and…
    • The canonical point $x$ of $\gamma$ has a path through the edges of $\gamma$ to $v$ whose intersection with the edges of $\pi$ has size $b_{1}+b_{2}$ modulo $2$ (i.e., if $b_{1}=b_{2}$, the size must be even, and if $b_{1}\neq b_{2}$, the size must be odd). Equivalently, the projected connected component of $\gamma$ containing the canonical point copy $(\gamma,b_{2})$ also contains $v^{b_{1}}$.
We visualize an example of the auxiliary graph in Figure 10.
(a) A canonical point is added along each obstacle, and edges are created from the canonical point to each arrangement vertex along the obstacle, subject to the parity conditions.
(b) A shortest-path between the two copies of an obstacle’s canonical point. The obstacles whose canonical points lie in this path form a set of obstacles separating $s$ and $t$.
(c) The set of obstacles induced by the shortest path in Figure 10(b).
Figure 10: The construction of the auxiliary graph for the homology cover given in Figure 9.
The auxiliary graph is useful because the shortest path between any two corresponding vertex copies or (separately) any two corresponding canonical point copies both correspond exactly to the solution to the $(s,t)$ point-separation problem. The high-level construction to prove this fact is as follows:
@@
(a) The intersection graph with the path $\pi$.
(b) The intersection graph after deleting the edges whose underlying paths cross $\pi$ an odd number of times.
(c) Two copies of the intersection graph are connected with “crossing” edges (in purple).
(d) A shortest-path in the homology cover between two copies of an obstacle.
Figure 11: How the intersection graph can be transformed into the intersection graph in the homology cover.
For the oracle model, the purpose of the oracles will be to construct the geometric intersection graph of the obstacle copies in the homology cover (which we will denote as $\overline{G}$). We will call $\overline{G}$ the intersection graph in the homology cover. See Figure 11 for an example. We will present two different constructions of this graph. They are equivalent, and each of them is quite simple, but they will serve two different purposes: The first will clarify which types of oracle queries are necessary, and provide an algorithmic construction. The second will instead build on the tools we have for the arrangement model, proving the correctness of another shortest-paths approach.
Let $\mathcal{C}$ denote the set of obstacles in the plane. We assume the canonical points for each obstacle are given (or implied) – they will be used by the oracle. We also assume some simple $s-t$ path $\pi$ is fixed – this will also be used by the oracle. In most cases, $\pi$ will be $\overline{st}$. The first construction of the graph is as follows:
@@
Testing for this condition is actually the only type of oracle query needed (although we will also require support for the case when $\gamma_{1}=\gamma_{2}$ later). Cabello and Giannopoulos [CabelloG16] used a larger set of query types, but together they can be used to perform this type by carefully choosing the canonical points and fixing $\pi$ to $\overline{st}$, so our variant of the oracle model is slightly more general (although practically equivalent).
@@
  • Therefore, the set of vertices visited by the shortest path from some $(\gamma,-)$ to $(\gamma,+)$ (over all possible $\gamma\in\mathcal{C}$, breaking ties arbitrarily) corresponds to a minimum-weight set of obstacles separating $s$ and $t$.